# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v1

@@ 1 Introduction
$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}})$ $:=(f_{1}({\bm{\theta}}),\cdots,f_{m}({\bm{\theta}}))^{\top},$ (1)
@@
Different from traditional single-objective optimization, which focuses on improving a single metric, MOO aims to obtain different solutions that achieve different trade-offs among objectives. Particularly, in MOO, users can specify their preferences regarding the relative importance of objectives, typically represented by a preference vector ${\bm{\alpha}}=[\alpha_{1},\cdots,\alpha_{m}]\in\Delta_{m-1}$, where each $\alpha_{i}$ indicates the importance of the $i$-th objective and $\Delta_{m-1}$ denotes a $(m-1)$-dimensional simplex. MOO is expected to achieve a solution that matches the user-defined preference. Hence, compared to single-objective optimization, MOO is more practical for complex real-world applications.
@@
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [Gradient-based Multi-objective Deep Learning, fill=gray!45, parent [Finding a Single Solution (§3), for tree=single [Loss Balancing Methods (§3.1), single [DWA [9]; UW [10]; IMTL-L [11]; Lin et al. [12]; MOML [13, 14]; Auto-$\lambda$ [15]; FORUM [16]; RW [17]; STCH [18], single_work] ] [Gradient Balancing Methods (§3.2), single [Gradient Weighting Methods (§3.2.1), single [MGDA [6]; CAGrad [19]; SDMGrad [20]; IMTL-G [11]; Nash-MTL [21]; FairGrad [22]; Aligned-MTL [23]; GradNorm [24]; DB-MTL [12], single_work] ] [Gradient Manipulation Methods (§3.2.2), single [PCGrad [25]; GradVac [26]; GradDrop [27], single_work] ] ] ] [Finding a Finite Set of Solutions (§4), for tree=finite [Preference Vector-based Methods (§4.1), finite [PMTL [28]; EPO [29, 30]; WC-MGDA [31]; PMGDA [32]; FERERO [33]; GMOOAR [34]; UMOD [35], finite_work] ] [Preference Vector-free Methods (§4.2), finite [HVGrad [36, 37, 38]; MOO-SVGD [39]; F4M [40]; SoM [41]; MosT [42], finite_work] ] ] [Finding an Infinite Set of Solutions (§5), for tree=infinite [Hypernetwork-based Methods (§5.1.1), infinite [PHN [43]; CPMTL [44]; PHN-HVI [45]; SUHNPF [46]; Tuan et al. [47], infinite_work] ] [Preference-Conditioned Network-based Methods (§5.1.2), infinite [COSMOS [48]; YOTO [49]; Raychaudhuri et al. [50], infinite_work] ] [Model Combination-based Methods (§5.1.3), infinite [PaMaL [51]; LORPMAN [52]; Chen and Kwok [53]; Panacea [54]; Tang et al. [55]; PaLoRA [56], infinite_work] ] ] ]
@@ 1.1 Taxonomy
Instead of the traditional approach of directly summing all objectives into a single scalar function, MOO algorithms in deep learning can be categorized into three main approaches based on the goal:
  • When the goal is to obtain a single, well-balanced solution, techniques like loss balancing and gradient balancing are employed to resolve conflicts between objectives during optimization. These methods focus on appropriately adjusting the contributions of each objective to achieve a solution that is not overly biased toward any one objective.
  • When the goal is to provide a finite set of Pareto optimal solutions, methods either divide the problem into subproblems using preference vectors or directly optimize for multiple solutions simultaneously. This allows users to select from a set of solutions that best meet their specific needs.
  • When the goal is to generate an infinite set of Pareto optimal solutions, the focus shifts to learning a subspace of solutions rather than individual ones. Algorithms in this category propose efficient structures to learn this subspace. During training, preference vectors are randomly sampled in each iteration. Once a preference is sampled, the optimization reduces to solving for either a single solution or a finite set, making it possible to use techniques from the first two approaches.
@@ 1.2 Comparison with Related Surveys
In contrast, this survey focuses on recent gradient-based MOO methods in deep learning, encompassing approaches for finding a single solution, a finite set of solutions, and an infinite set of solutions, and covering algorithms, theories, and applications. To the best of our knowledge, it is the first survey paper to focus on the gradient-based MOO methods in deep learning.
@@ 1.3 Organization of the Survey
The notations used in this survey are summarized in Table I. The rest of this survey is organized as follows:
@@
TABLE I: Summary of Notations.
Notation Description
$m$ Number of objectives.
$n$ Number of finite Pareto solutions.
$K$ Number of iterations.
${\bm{f}}=[f_{1},\cdots,f_{m}]$ Objective function.
${\bm{\lambda}}=[\lambda_{1},\cdots,\lambda_{m}]$ Objective weight vector.
${\bm{\alpha}}=[\alpha_{1},\cdots,\alpha_{m}]$ Preference vector.
${\bm{z}}^{*}=[z_{1}^{*},\cdots,z_{m}^{*}]$ Ideal objective value.
${\bm{\theta}}^{(k)}$ Solution at $k$-th iteration.
${\bm{\theta}}^{(1)},\cdots,{\bm{\theta}}^{(n)}$ Solution $1$ to Solution $n$ in a size-$n$ solution set.
${\bm{g}}^{(k)}_{i}$ The gradient vector of $i$-th objective at $k$-th iteration.
${\bm{G}}^{(k)}=[{\bm{g}}_{1}^{(k)},\cdots,{\bm{g}}_{m}^{(k)}]$ Gradient matrix at $k$-th iteration.
${\bm{d}}^{(k)}$ Updating direction of ${\bm{\theta}}$ at $k$-th iteration.
${\bm{\phi}}$ Parameters of Pareto set learning structures.
$[m]$ Index set $\{1,\cdots,m\}$.
[…]
$\eta$ Step size for updating ${\bm{\theta}}$.
$\epsilon$ Error tolerance.
@@ 2 Preliminary: Key Concepts in MOO
In this section, we first introduce some Pareto concepts in Section 2.1 and then review two classical methods (i.e., linear scalarization and Tchebycheff scalarization) to solve MOO problems in Section 2.2.
@@ 2.1 Pareto Concepts
Definition 3 (Pareto set (PS) [1])
A PS is the set of all Pareto optimal solutions.
Definition 4 (Pareto front (PF) [1])
A PF is the set of all objective function values of the PO solutions.
Figure 2 illustrates the aforementioned definitions. The blue circles represent Pareto optimal solutions, the yellow circles represent weakly Pareto optimal solutions, and the red circles are solutions dominated by at least one other solution. The blue curve represents the Pareto front.
@@
In deep learning, the parameter ${\bm{f}}({\bm{\theta}})$ is typically optimized using gradient descent. In single-objective optimization, when the objective function is non-convex, the optimization process often reaches a stationary point. In the case of MOO, the solution generally converges to a Pareto stationary solution, which is defined as follows:
@@
Given a solution set ${\mathbb{S}}=\{{\bm{q}}^{(1)},\cdots,{\bm{q}}^{(N)}\}$ and a reference point ${\bm{r}}$, the hypervolume of ${\mathbb{S}}$ is calculated by:
@@
HV provides a unary measure of both convergence and diversity of objective vectors. A larger HV indicates that the obtained solutions are diverse and their objective values are close to the PF.
@@ 2.2 Scalarization Methods
Linear scalarization. The most straightforward approach to solving MOO problems (i.e., problem (1)) is to reformulate them as single-objective optimization problems weighted by the given preference vector ${\bm{\alpha}}$ as follows:
@@
Linear scalarization is widely used in practice due to its simplicity. However, this method suffers from a significant limitations: It can only identify solutions on the convex part of the PF. For PFs with a concave shape, it typically finds only the endpoints.
Tchebycheff scalarization. Another popular scalarization function is Tchebycheff scalarization [66, 67], which is defined as:
@@
where $z_{i}^{*}$ are reference values, which can be set as the ideal value of the $i$-th objective. Unlike linear scalarization, Tchebycheff scalarization can explore the entire PF by traversing all preference vectors within the preference simplex. This ensures that both convex and concave regions of the PF are covered. The relationship between Tchebycheff scalarization and weak Pareto optimality can be formally described as follows:
@@
A solution ${\bm{\theta}}^{*}$ is weakly Pareto optimal of the original MOO problem (problem (1)) if and only if there exists a preference vector ${\bm{\alpha}}$ such that ${\bm{\theta}}^{*}$ is an optimal solution of problem (4).
If ${\bm{\theta}}^{*}$ is unique for a given ${\bm{\alpha}}$, it is considered Pareto optimal. However, Tchebycheff scalarization poses challenges for gradient-based deep learning due to the nonsmooth nature of the $\max(\cdot)$ operator. This nonsmoothness leads to non-differentiability and a slow convergence rate of $\mathcal{O}(1/\epsilon^{2})$, where $\epsilon$ represents the error tolerance [18]. Even when all objectives $\{f_{i}\}_{i=1}^{m}$ are differentiable, the problem (4) remains non-differentiable.
@@ 3 Finding a single Pareto optimal solution
In some scenarios, it is sufficient to identify a single Pareto optimal solution without the need to explore the entire Pareto set. Here, all objectives are treated as equally important (i.e., the preference vector is ${\bm{\alpha}}=[\frac{1}{m},\cdots,\frac{1}{m}]$), and the goal is to produce a single model that effectively balances all objectives. Most approaches in multi-task learning (MTL) [3, 69, 57] belong to this category.
The most straightforward method is the linear scalarization method (i.e., problem (3)) with $\alpha_{i}=\frac{1}{m}$, which is known as Equal Weighting (EW) [3] in MTL. However, EW may cause some tasks to have unsatisfactory performance [70, 12]. Therefore, to improve the performance, many methods have been proposed to dynamically tune the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training, which can in general be formulated as:
@@ 3.1 Loss Balancing Methods
Dynamic Weight Average (DWA) [9] estimates each objective weight as the ratio of the training losses in the last two iterations for the corresponding objective. Formally, the objective weight in $k$-th iteration ($k>2$) is computed as follows:
@@
where ${\bm{s}}$ is learnable and $\log s_{i}$ is a regularization term.
@@
Ye et al. [13, 14] propose Multi-Objective Meta Learning (MOML) to adaptively tune the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ according to the validation performance by reformulating problem (5) as a multi-objective bi-level optimization problem:
$\min_{{\bm{\lambda}}}~{}\big{(}f_{1}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{1}),\cdots,f_{m}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{m})\big{)}^{\top}$ (9a)
@@
where $\mathcal{D}^{\text{tr}}_{i}$ and $\mathcal{D}^{\text{val}}_{i}$ are the training and validation datasets for the $i$-th objective, respectively. In each training iteration, when given objective weights $\{\lambda_{i}\}_{i=1}^{m}$, MOML first learns the model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the training dataset by solving the lower-level (LL) subproblem (9b) with $T$ iterations and then updates ${\bm{\lambda}}$ in the upper-level (UL) subproblem (9a) via minimizing the loss of the trained MTL model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the validation dataset of each objective. However, when solving the UL subproblem (9a), MOML needs to calculate the complex hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ which requires to compute many Hessian-vector products via the chain rule. Hence, the time and memory costs of MOML grow significantly fast with respect to the dimension of ${\bm{\theta}}$ and the number of LL iterations $T$.
Auto-$\lambda$ [15] has a similar formulation with problem (9) by replacing the multi-objective UL subproblem (9a) with a single-objective problem (i.e., $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}f_{i}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{i})$). Besides, Auto-$\lambda$ approximates the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ via the finite difference approach. Thus, it is more efficient than MOML [13, 14].
Recently, FORUM [16] is proposed to improve the efficiency of MOML [13, 14]. It first reformulates the problem (9) as a constrained MOO problem via the value-function approach [71] and then proposes a multi-gradient aggregation method to solve it. Compared with MOML, FORUM is a first-order method and does not need to compute the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$. Thus, it is significantly more efficient than MOML.
@@
Different from the above methods which are based on linear scalarization, Lin et al. [18] propose the Smooth Tchebycheff scalarization (STCH) to make the Tchebycheff scalarization a smooth function, which is defined as:
$\min_{{\bm{\theta}}}f_{\bm{\alpha}}^{\mathrm{STche}}({\bm{\theta}},\mu):=\mu\log\left(\sum_{i=1}^{m}\exp\left(\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*})/\mu\right)\right),$ (10)
where $\mu>0$ is a smoothing parameter, and ${\bm{z}}^{*}$ is a reference point which is the same as in the Tchebycheff scalarization (problem (4)). STCH improves upon the original Tchebycheff scalarization by making $f_{\bm{\alpha}}^{\mathrm{STche}}({\bm{\theta}},\mu)$ smooth when all objective functions $f_{i}({\bm{\theta}})$ are smooth, resulting in faster convergence. It also retains the advantages of the original Tchebycheff scalarization: (1) convexity when all objective functions are convex and (2) the Pareto optimality property by appropriately choosing $\mu$.

### v2

@@ 1 Introduction
$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}})$ $:=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$ (1)
@@
Figure 1: An overview of multi-objective deep learning.
Different from traditional single-objective optimization, which focuses on improving a single metric, MOO aims to obtain different solutions that achieve different trade-offs among objectives. Particularly, in MOO, users can specify their preferences regarding the relative importance of objectives, typically represented by a preference vector ${\bm{\alpha}}=[\alpha_{1},\dots,\alpha_{m}]^{\top}\in\Delta_{m-1}$, where each $\alpha_{i}$ indicates the importance of the $i$-th objective and $\Delta_{m-1}$ denotes a $(m-1)$-dimensional simplex. MOO is expected to achieve a solution that matches the user-defined preference. Hence, compared to single-objective optimization, MOO is more practical for complex real-world applications.
@@
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [ Gradient-based Multi-objective Deep Learning , fill=gray!45, parent [Finding a Single Solution (§3), for tree=single [Loss Balancing Methods (§3.1), single [DWA [9]; UW [10]; IMTL-L [11]; Lin et al. [12]; MOML [13, 14]; Auto-$\lambda$ [15]; FORUM [16]; RW [17]; STCH [18], single_work] ] [Gradient Balancing Methods (§3.2), single [Gradient Weighting Methods (§3.2.1), single [MGDA [6]; CAGrad [19]; SMG [20]; CR-MOGM [21]; MoCo [22]; PSMGD [23]; MoDo [24]; SDMGrad [25]; SGSMGrad [26]; IMTL-G [11]; Nash-MTL [27]; FairGrad [28]; UPGrad [29]; Aligned-MTL [30]; GradNorm [31]; DB-MTL [12]; FAMO [32], single_work] ] [Gradient Manipulation Methods (§3.2.2), single [PCGrad [33]; GradVac [34]; GradDrop [35], single_work] ] ] ] [Finding a Finite Set of Solutions (§4), for tree=finite [Methods Based on Preference Vectors (§4.1), finite [PMTL [36]; EPO [37, 38]; WC-MGDA [39]; PMGDA [40]; FERERO [41]; GMOOAR [42]; UMOD [43], finite_work] ] [Methods without Using Preference Vectors (§4.2), finite [GradHV [44, 45, 46, 47]; MOO-SVGD [48]; F4M [49]; SoM [50]; MosT [51], finite_work] ] ] [Finding an Infinite Set of Solutions (§5), for tree=infinite [Hypernetwork-based Methods (§5.1.1), infinite [PHN [52]; CPMTL [53]; PHN-HVI [54]; SUHNPF [55]; Hyper-Trans [56], infinite_work] ] [Preference-Conditioned Network-based Methods (§5.1.2), infinite [COSMOS [57]; YOTO [58]; Raychaudhuri et al. [59], infinite_work] ] [Model Combination-based Methods (§5.1.3), infinite [PaMaL [60]; LORPMAN [61]; Chen and Kwok [62]; Panacea [63]; Tang et al. [64]; PaLoRA [65], infinite_work] ] ] ]
@@ 1.1 Taxonomy
Instead of the traditional approach of directly summing all objectives into a single scalar function, MOO algorithms in deep learning can be categorized into three main approaches:
  • To obtain a single, well-balanced solution, techniques like loss balancing and gradient balancing are employed to resolve conflicts between objectives during optimization. These methods focus on adjusting the weight of each objective to achieve a solution that is not overly biased toward any one objective.
  • To obtain a finite set of Pareto optimal solutions, methods either divide the problem into subproblems using preference vectors or directly optimize for multiple solutions simultaneously. This allows users to select from a set of solutions that best meet their specific needs.
  • To generate an infinite set of Pareto optimal solutions, the focus shifts to learning a subspace of solutions rather than individual ones. Algorithms in this category propose efficient structures to learn this subspace. During training, preference vectors are randomly sampled in each iteration. Once a preference is sampled, the optimization reduces to solving for either a single solution or a finite set, making it possible to use techniques from the first two approaches.
@@ 1.2 Comparison with Related Surveys
Different with the previous surveys, this survey focuses on recent gradient-based MOO methods in deep learning, encompassing approaches for finding a single solution, a finite set of solutions, and an infinite set of solutions, and covering algorithms, theories, and applications. To the best of our knowledge, it is the first survey paper to focus on the gradient-based MOO methods in deep learning.
TABLE I: Summary of Notations.
Notation Description
$m$ Number of objectives.
$n$ Number of finite Pareto solutions.
$K$ Number of iterations.
${\bm{f}}=[f_{1},\dots,f_{m}]^{\top}$ Objective function.
${\bm{\lambda}}=[\lambda_{1},\dots,\lambda_{m}]^{\top}$ Objective weight vector.
${\bm{\alpha}}=[\alpha_{1},\dots,\alpha_{m}]^{\top}$ Preference vector.
${\bm{z}}^{*}=[z_{1}^{*},\dots,z_{m}^{*}]^{\top}$ Ideal objective value.
${\bm{\theta}}^{(k)}$ Solution at $k$-th iteration.
${\bm{\theta}}^{(1)},\dots,{\bm{\theta}}^{(n)}$ Solution $1$ to Solution $n$ in a size-$n$ solution set.
${\bm{g}}^{(k)}_{i}$ The gradient vector of $i$-th objective at $k$-th iteration.
${\bm{G}}^{(k)}=[{\bm{g}}_{1}^{(k)},\dots,{\bm{g}}_{m}^{(k)}]$ Gradient matrix at $k$-th iteration.
${\bm{d}}^{(k)}$ Updating direction of ${\bm{\theta}}$ at $k$-th iteration.
${\bm{\phi}}$ Parameters of Pareto set learning structures.
$[m]$ Index set $\{1,\dots,m\}$.
[…]
$\eta$ Step size for updating ${\bm{\theta}}$.
$\epsilon$ Error tolerance.
@@ 1.3 Organization of the Survey
The notations used in this survey are summarized in Table I. An overview of MOO in deep learning is provided in Figure 1. The rest of this survey is organized as follows:
@@ 2 Preliminary: Key Concepts in MOO
In this section, we first introduce some Pareto concepts in Section 2.1 and then review two classical methods (i.e., linear scalarization and Tchebycheff scalarization) to find Pareto solutions of MOO problems in defined in Section 2.2. An illustration of some Pareto concepts is in Appendix A.
@@ 2.1 Pareto Concepts
Definition 3 (Pareto set (PS) and Pareto front (PF) [1])
A PS is the set of all PO solutions. A PF is the set of all objective function values of the PO solutions.
@@
In deep learning, the parameter ${\bm{\theta}}$ is typically optimized using gradient descent. In single-objective optimization, when the objective function is non-convex, the optimization process often reaches a stationary point. In the case of MOO, the solution generally converges to a Pareto stationary solution, which is defined as follows:
@@
Given a solution set ${\mathbb{S}}=\{{\bm{q}}^{(1)},\dots,{\bm{q}}^{(N)}\}$ and a reference point ${\bm{r}}$, the hypervolume of ${\mathbb{S}}$ is calculated by:
@@
HV provides a unary measure of both convergence and diversity of a solution set. A larger HV indicates that the obtained solutions are diverse and their objective values are close to the PF.
@@ 2.2 Scalarization Methods
Linear scalarization. The most straightforward way to solve MOO problems is to reformulate them as single-objective optimization problems weighted by the given preference vector ${\bm{\alpha}}$ as follows:
@@
Linear scalarization is widely used in practice due to its simplicity. However, this method suffers from a limitation: it can only identify solutions on the convex part of the PF. For PFs with a concave shape, using LS only finds the endpoints of a PF. Tchebycheff scalarization. Another way to convert an MOP to a single one is to use the Tchebycheff scalarization function [75, 76], defined as:
@@
where $z_{i}^{*}$ are reference values, usually set as the ideal value of the $i$-th objective. Unlike linear scalarization, Tchebycheff scalarization can explore the entire PF by traversing all preference vectors within the simplex. This ensures that both convex and concave regions of the PF are covered. The relationship between Tchebycheff scalarization and weak Pareto optimality can be formally described as follows:
@@
A solution ${\bm{\theta}}^{*}$ of the original MOO problem (1) is weakly Pareto optimal if and only if there exists a preference vector ${\bm{\alpha}}$ such that ${\bm{\theta}}^{*}$ is an optimal solution of problem (4).
If ${\bm{\theta}}^{*}$ is unique for a given ${\bm{\alpha}}$, it is considered Pareto optimal. However, Tchebycheff scalarization poses challenges for gradient-based deep learning due to the nonsmooth nature of the $\max(\cdot)$ operator. This nonsmoothness leads to non-differentiability and a slow convergence rate of $\mathcal{O}(1/\epsilon^{2})$, where $\epsilon$ represents the error tolerance [18]. Even when all objectives $\{f_{i}\}_{i=1}^{m}$ are differentiable, problem (4) remains non-differentiable.
@@ 3 Finding a Single Pareto Optimal Solution
In some scenarios such as multi-task learning (MTL) [3, 78, 66], it is sufficient to identify a single Pareto optimal solution without the need to explore the entire Pareto set. Here, all objectives are usually treated as equally important (i.e., the preference vector is ${\bm{\alpha}}=[\frac{1}{m},\dots,\frac{1}{m}]^{\top}$), and the goal is to produce a single model that effectively balances all objectives. The most straightforward method is the linear scalarization method (i.e., problem (3)) with $\alpha_{i}=\frac{1}{m}$, which is known as Equal Weighting (EW) [3] in MTL. However, EW may cause some tasks to have unsatisfactory performance [79, 12]. Therefore, to improve the performance, many methods have been proposed to dynamically tune the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training, which can in general be formulated as:
@@ 3.1 Loss Balancing Methods
Dynamic Weight Average (DWA) [9] estimates each objective weight as the ratio of the training losses in the last two iterations. Formally, the objective weight in $k$-th iteration ($k>2$) is computed as follows:
@@
where ${\bm{s}}=[s_{1},\dots,s_{m}]^{\top}$ is learnable and $\log s_{i}$’s are regularization terms.
@@
Ye et al. [13, 14] propose Multi-Objective Meta Learning (MOML), which adaptively tunes the objective weights ${\bm{\lambda}}$ based on the validation performance, reformulating problem (5) as a multi-objective bi-level optimization problem:
$\min_{{\bm{\lambda}}}~{}\big{[}f_{1}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{1}),\dots,f_{m}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{m})\big{]}^{\top}$ (9a)
@@
where $\mathcal{D}^{\text{tr}}_{i}$ and $\mathcal{D}^{\text{val}}_{i}$ are the training and validation datasets for the $i$-th objective, respectively. In each training iteration, when given objective weights $\{\lambda_{i}\}_{i=1}^{m}$, MOML first learns the model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the training dataset by solving lower-level (LL) subproblem (9b) with $T$ iterations and then updates ${\bm{\lambda}}$ in the upper-level (UL) subproblem (9a) via minimizing the loss of the trained MTL model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the validation dataset of each objective. However, when solving the UL subproblem (9a), MOML needs to calculate the complex hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ which requires to compute many Hessian-vector products via the chain rule. Hence, the time and memory costs of MOML grow significantly fast with respect to the dimension of ${\bm{\theta}}$ and the number of LL iterations $T$.
Auto-$\lambda$ [15] modifies problem (9) by replacing the multi-objective UL subproblem (9a) with a single-objective problem: $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}f_{i}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{i})$. Besides, Auto-$\lambda$ approximates the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ via the finite difference approach. Thus, it is more efficient than MOML [13, 14].
Recently, FORUM [16] is proposed to improve the efficiency of MOML [13, 14]. FORUM first reformulates problem (9) as a constrained MOO problem via the value-function approach [80] and then proposes a multi-gradient aggregation method to solve it. Compared with MOML, FORUM is a first-order method and does not need to compute the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$. Thus, it is significantly more efficient than MOML.
@@
Different from the above methods which are based on linear scalarization, Lin et al. [18] propose Smooth Tchebycheff scalarization (STCH), which transforms the Tchebycheff scalarization into a smooth function:
$\min_{{\bm{\theta}}}f_{\bm{\alpha}}^{\mathrm{STCH}}({\bm{\theta}},\mu):=\mu\log\sum_{i=1}^{m}\exp\left(\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*})/\mu\right),$ (10)
where $\mu>0$ is a smoothing parameter, and ${\bm{z}}^{*}$ is a reference point as in Tchebycheff scalarization (problem (4)). STCH improves upon the original Tchebycheff scalarization by making $f_{\bm{\alpha}}^{\mathrm{STCH}}({\bm{\theta}},\mu)$ smooth when all objective functions $f_{i}({\bm{\theta}})$’s are smooth, resulting in faster convergence. It also retains the advantages of the original Tchebycheff scalarization: (1) convexity when all objective functions are convex and (2) Pareto optimality by appropriately choosing $\mu$.